OpenStudy (anonymous):

fifteen times the square of a non-zero number is equal to 20 times the number

OpenStudy (anonymous):

Helpp Pleasee

OpenStudy (harsimran_hs4):

let the number be x square of the number is : x^2 so the equation becomes 15*(x^2) = 20*x can you solve it now ?

OpenStudy (anonymous):

Noo..(: Answerr Please

OpenStudy (harsimran_hs4):

15x^2 - 20x = 0 5x(3x - 4) = 0 tell me the solutions now ?

OpenStudy (anonymous):

I Need dha answer im doing a test lol ian got time to be thinking and solving

OpenStudy (harsimran_hs4):

now you should be able to Here we guide and help and WE DON`T give direct answer i have taken you to last step i want you to atleast try

OpenStudy (anonymous):

I like @harsimran_hs4 's approach. That's fair.

OpenStudy (anonymous):

Do i Multiply the 5 and 3 ?? and then subtract 4?

OpenStudy (harsimran_hs4):

consider pq = 0 , so what are the possible solutions either p = 0 or q = 0 or both are zero now if you carefully observe in above equation p = 5x and q = 3x-4 now tell what are possible values of x ?

OpenStudy (anonymous):

its 5x and 3x ???

OpenStudy (harsimran_hs4):

look either 5x can be zero or 3x-4 can be zero or both if 5x = 0 then x = 0 if 3x-4 = 0 then x = 4/3 but the number is non zero so x = 4/3

OpenStudy (harsimran_hs4):

since you had problem understanding , try and analyze each step and ask if there is any doubt

OpenStudy (anonymous):

It Says That Was Wrong But Thanks Anyway (:

OpenStudy (anonymous):

Really? \(\frac{4}{3}\) Is wrong? I don't think so... \(15(\frac{4}{3})^2=20\frac{4}{3}\) \(15(\frac{16}{9})=\frac{80}{3}\) \(\frac{240}{9}=\frac{80}{3}\)

OpenStudy (harsimran_hs4):

and yes since it was a quadratic eq it can have atmost 2 real solutions and they are 0 (which you can easily see) and 4/3 (which @KeithAfasCalcLover showed above) so check if you copied the statement correctly

OpenStudy (anonymous):

k